[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Tue Nov 15, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Tue Nov 15 18:03:10 UTC 2016


HSU Shift Log for Tue Nov 15 2016 13:03:03 GMT-0500 (EST)

Shift: H

Forecaster: SR Stewart

Technical Issues Summary:

ATCF issues: Backup ATCF was running slow when trying to issue TWOAT  
yesterday and this morning at 12Z due to dialogue boxes hanging up for  
about a minute or so before TWO was finally transmitted; rebooted ATCF and  
system now appears to be functioning properly.

N-AWIPS issues: Recently rebuilt RIDGES N-AWIPS system on Atlantic desk  
still has some issues with the dialogue boxes having the 'jitters' and a  
light blue color band about 2 inches wide extending across the bottom of  
the upper monitor screen; TSB was notified and they repaired a loose cable  
that was causing the problem.

AWIPS-2/NCP issues: None.

Other Techincal issues: Conducted P-Surge communications test on AL852016  
for SDM and TSB at 1432Z; called SDM at 1502Z and he confirmed that the  
P-Surge test had been successful.

Miscellaneous issues: None.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20161115/b44163ed/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list